# On Vanishing Gradients, Over-Smoothing, and Over-Squashing in GNNs: Bridging Recurrent and Graph Learning

### v1

@@ Abstract
Graph Neural Networks (GNNs) are models that leverage the graph structure to transmit information between nodes, typically through the message-passing operation. While widely successful, this approach is well-known to suffer from the over-smoothing and over-squashing phenomena, which result in representational collapse as the number of layers increases and insensitivity to the information contained at distant and poorly connected nodes, respectively. In this paper, we present a unified view of these problems through the lens of vanishing gradients, using ideas from linear control theory for our analysis. We propose an interpretation of GNNs as recurrent models and empirically demonstrate that a simple state-space formulation of a GNN effectively alleviates over-smoothing and over-squashing at no extra trainable parameter cost. Further, we show theoretically and empirically that (i) GNNs are by design prone to extreme gradient vanishing even after a few layers; (ii) Over-smoothing is directly related to the mechanism causing vanishing gradients; (iii) Over-squashing is most easily alleviated by a combination of graph rewiring and vanishing gradient mitigation. We believe our work will help bridge the gap between the recurrent and graph neural network literature and will unlock the design of new deep and performant GNNs.
Machine Learning, ICML
@@ 1 Introduction
Graph Neural Networks (GNNs) (Sperduti, 1993; Gori et al., 2005; Scarselli et al., 2008; Bruna et al., 2014; Defferrard et al., 2017) have become a widely used architecture for processing information on graph domains. Most GNNs operate via message passing, where information is exchanged between neighboring nodes, giving rise to Message-Passing Neural Networks (MPNNs). Some of the most popular instances of this type of architecture include GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2018), and GraphSAGE (Hamilton et al., 2017).
Despite its widespread use, this paradigm also suffers from some fundamental limitations. Most importantly, we highlight the issue of over-smoothing (Nt & Maehara, 2019; Cai & Wang, 2020; Rusch et al., 2023a), where feature representations become exponentially similar as the number of layers increases, and over-squashing (Alon & Yahav, 2021; Topping et al., 2021; Di Giovanni et al., 2023), which describes the difficulty of propagating information across faraway nodes, as the exponential growth in a node’s receptive field results in many messages being compressed into fixed-size vectors. Although these two issues have been studied extensively, and there exists evidence that they are trade-offs of each other (Giraldo et al., 2023), there is no unified theoretical framework that explains why architectures which solve these problems work and whether there exists a common underlying cause that governs these problems.
In this work, we analyze over-smoothing and over-squashing from the lens of vanishing gradients. In particular, we ask several questions about the appearance and consequences of this phenomenon in GNNs: (i) How prone are GNNs to gradient vanishing? (ii) What is the effect of gradient vanishing on over-smoothing? (iii) Can preventing vanishing gradients effectively mitigate over-squashing? (iv) Can methods used in the non-linear (Hochreiter & Schmidhuber, 1997; Pascanu et al., 2013; Beck et al., 2024) and more recently linear (Gu & Dao, 2023; Orvieto et al., 2023) recurrent neural network (RNN) literature be effective at dealing with over-smoothing and over-squashing? With the latter questions, we aim to fill the gaps and open questions left in the over-squashing analysis of Di Giovanni et al. (2023). Further, we hope the point on over-smoothing will help provide a theoretical explanation for why certain architectural choices have led to the design of deep and performant GNNs.
@@
  • In Section 3, we explore the connection between GNNs and sequence models and demonstrate how graph convolutional and attentional models are susceptible to a phenomenon we term extreme gradient vanishing. We propose GNN-SSM, a GNN model that is written as a state-space model, allowing for a better control of the spectrum of the Jacobian.
  • In Section 4, we show how vanishing gradients contribute to over-smoothing, providing a much more precise explanation for why over-smoothing occurs in practice in GNNs by studying the spectrum of the layer-wise Jacobians. We show that GNN-SSMs are able to exactly control the rate of smoothing.
@@ 2.1 Message Passing Neural Networks
for $k=\{1,\ldots,K\}$, where $\psi^{(k)}$ is a permutation invariant aggregation function and $\phi^{(k)}$ combines the incoming messages from one’s neighbors with the previous embedding of oneself to produce an updated representation. The most commonly used aggregation function takes the form
@@
where $\tilde{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, and $\mathbf{D}\in\mathbb{R}^{n\times n}$ is a diagonal matrix where $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. One can also consider a matrix representation of the features $\mathbf{H}^{(k)}\in\mathbb{R}^{n\times d_{k}}$. Throughout the paper, we will use the terms GNN and MPNN interchangeably, and will generally consider the most widely used instance of GNNs, which are Graph Convolutional Networks (GCNs) (Kipf & Welling, 2017) whose matrix update equation is given by:
@@
where $\hat{\mathbf{A}}=\left(\mathbf{D}+\mathbf{I}\right)^{-1/2}\left(\mathbf{A}+\mathbf{I}\right)\left(\mathbf{D}+\mathbf{I}\right)^{-1/2}$ is the adjacency matrix with added self connections through the identity matrix $\mathbf{I}$, and $\sigma(\cdot)$ is a nonlinearity. Our analysis also applies to Graph Attention Networks (GATs) (Veličković et al., 2018), where the fixed normalized adjacency is replaced by a learned adjacency matrix which dynamically modulates connectivity while preserving the key spectral properties used in our analysis.
Figure 1: Latent evolution of 2-dimensional node features when passing through layers of a GNN-SSM with $\text{eig}(\Lambda)\approx 1$. The colors indicate the norm of the feature at each node, and the vector field indicates direction.
@@ 2.2 Recurrent Neural Networks
This type of approach has deep connections with ideas from dynamical systems (Strogatz, 2018) and chaotic systems (Engelken & Wolf, 2023). These ideas have become more relevant in recent work (Gu et al., 2019; Orvieto et al., 2023), where the nonlinearity in (4) is removed in the interest of parallelization and the ability to directly control the dynamics of the system through the eigenvalues of state transition matrix $\mathbf{W}_{h}$. We note that these types of approaches are also popular in the reservoir computing literature (Jaeger, 2001), where the state transition matrix is left untrained and more emphasis is placed on the dynamics of the model.
@@ 2.3 The Vanishing and Exploding Gradient Problem
$\frac{\partial\mathcal{L}}{\partial\mathbf{\boldsymbol{\theta}}^{(i)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{(K)}}\prod_{k=i+1}^{K}\frac{\partial\mathbf{H}^{(k)}}{\partial\mathbf{H}^{(k-1)}}\frac{\partial\mathbf{H}^{(i)}}{\partial\mathbf{\boldsymbol{\theta}}^{(i)}},$ (5)
where matrix $\mathbf{H}^{(k)}$ in an RNN will contain a single state vector. As identified by Pascanu et al. (2013), a major issue in training this type of models arises from the product Jacobian, given by:
@@
In general, we have that if $||\mathbf{J}_{k}||_{2}\approx\lambda$ for all layers then $||\mathbf{J}||_{2}\leq\lambda^{K-i}$. This means that we require $\lambda\approx 1$ for gradients to neither explode nor vanish, a condition also known as edge of chaos.
@@ 2.4 Over-smoothing, Over-squashing, and Vanishing Gradients in GNNs
Over-smoothing.
Over-smoothing (Cai & Wang, 2020; Oono & Suzuki, 2020; Rusch et al., 2023a) describes the tendency of GNNs to produce smoother representations as more and more layers are added. In particular, this has been related to the convergence to the $1$-dimensional kernel of the graph Laplacian and equivalently as a minimization process of the Dirichlet energy (Di Giovanni et al., 2022). In Section 4, we study this issue from the lens of vanishing gradients and show that over-smoothing has a much more simple explanation: it occurs due to the norm-contracting nature of GNN updates.
@@
Over-squashing (Alon & Yahav, 2021; Topping et al., 2021; Di Giovanni et al., 2023; Barbero et al., 2024) was originally introduced as a bottleneck resulting from ‘squashing’ into node representations amounts of information that are growing potentially exponentially quickly due to the topology of the graph. It is often characterized by the quantity $\left\lVert\partial\mathbf{h}_{u}^{(K)}/\partial\mathbf{h}_{v}^{(0)}\right\rVert$ being low, implying that the final representation of node $u$ is not very sensitive to the initial representation at some other node $v$. While the relationship between over-squashing and vanishing gradients was hinted at by Di Giovanni et al. (2023), in Section 5 we explore this relationship in detail by showing that techniques aimed to mitigate vanishing gradients in sequence models help to mitigate over-squashing in GNNs.
@@
Vanishing gradients have been extensively studied in RNNs (Bengio, 1994; Hochreiter & Schmidhuber, 1997; Pascanu et al., 2013), while this problem has been surprisingly mostly overlooked in the GNN community. For a detailed discussion on the relevant literature, we point the reader to the Appendix E. We simply highlight that there are works that have seen success in taking ideas from sequence modelling (Rusch et al., 2022; Gravina et al., 2023; Rusch et al., 2023b; Wang et al., 2024; Behrouz & Hashemi, 2024; Kiani et al., 2024) or signal propagation (Epping et al., 2024; Scholkemper et al., 2024) and bridging them to GNNs, but they rarely have a detailed discussion on vanishing gradients. In Section 5, we show that vanishing gradient mitigation techniques from RNNs seem to be very effective towards the mitigation of over-smoothing and over-squashing in GNNs and argue that the two communities have at a fundamental level very aligned problems and goals.
@@ 3.1 Similarities and differences between learning on sequences and graphs
The GNN architectures that first popularized deep learning on graphs (Bruna et al., 2014; Defferrard et al., 2017) were initially presented as a generalization of Convolutional Neural Networks (CNNs) to irregular domains. GCNs (Kipf & Welling, 2017) subsequently restricted the architecture in Defferrard et al. (2017) to a one-hop neighborhood. While this is still termed “convolutional” (due to weight sharing across nodes), the iterative process of aggregating information from each node’s neighborhood can also be viewed as recurrent-like state updates.
If we consider an RNN unrolled over time, it forms a directed path graph feeding into a state node with a self-connection— making it a special case of a GNN. Conversely, node representations in GNNs can be stacked using matrix vectorization, allowing us to interpret GNN layer operations as iterative state updates. This connection suggests that the main difficulty faced by RNNs — namely the vanishing and exploding gradients problem (Pascanu et al., 2013) — may likewise hinder the learning ability of GNNs. We note, however, that one key difference between RNNs and GNNs is that RNN memory only depends on how much information is dissipated by the model during the hidden state update, whereas GNNs normalize messages by the inverse node degree, which introduces an additional information dissipation step that we will explore in more detail in Section 5.
@@ 3.2 Graph convolutional and attentional models are prone to extreme gradient vanishing
Lemma 3.1 (Spectrum of the Jacobian’s singular values).
@@
$\bigl{\{}\,\lambda_{i}^{2}\,\mu_{j}\;\bigm{|}\;i=1,\ldots,n,\;\;j=1,\ldots,d_{k}\bigr{\}}.\$
@@
$\mathbb{E}\bigl{[}\gamma_{i,j}\bigr{]}$ $=\lambda_{i}^{2}\,\sigma^{2},$ (7)
@@
Theorem 3.2 shows that the singular-value spectrum of the Jacobian is modulated by the squared spectrum of the normalized adjacency. Since $|\lambda_{i}|\leq 1$ for all eigenvalues of the normalized adjacency, the ability of GCNs to propagate gradients is strictly worse than that of RNNs or MLPs. In particular, iterating these operations causes the majority of the spectrum to shrink to zero more quickly than in classical deep linear (Saxe et al., 2013) or nonlinear (Pennington et al., 2017) networks. Moreover, using sigmoidal activations and orthogonal weights will not push the singular-value spectrum to the edge of chaos as in Pennington et al. (2017), due to the additional contraction from the adjacency. The effect of each operation on the layer-wise Jacobian is empirically demonstrated in Figure 2, which also showcases the contraction effect of the normalized adjacency. The figure reveals that even a single layer’s Jacobian exhibits a long tail of squared singular values near zero. This spectral structure leads to ill-conditioned gradient propagation and non-isometric (not norm-preserving) signal dynamics. The same results hold for GATs, as the adjacency still exhibits a contractive spectral structure despite being learned during training Finally, we also plot the Jacobian’s singular value spectrum for several GNN architectures in Appendix D.1, which shows how this phenomenon is also present in other GNNs, even though it is less pronounced.
@@
Figure 2: Histogram of eigenvalue modulus of the Jacobian for linear, linear convolutional, and nonlinear convolutional layers.

### v2

@@ Abstract
Graph Neural Networks (GNNs) are models that leverage the graph structure to transmit information between nodes, typically through the message-passing operation. While widely successful, this approach is well-known to suffer from representational collapse as the number of layers increases and insensitivity to the information contained at distant and poorly connected nodes. In this paper, we present a unified view of the appearance of these issues through the lens of vanishing gradients, using ideas from linear control theory for our analysis. We propose an interpretation of GNNs as recurrent models and empirically demonstrate that a simple state-space formulation of a GNN effectively alleviates these issues at no extra trainable parameter cost. Further, we show theoretically and empirically that (i) Traditional GNNs are by design prone to extreme gradient vanishing even after a few layers; (ii) Feature collapse is directly related to the mechanism causing vanishing gradients; (iii) Long-range modeling is most easily achieved by a combination of graph rewiring and vanishing gradient mitigation. We believe our work will help bridge the gap between the recurrent and graph learning literature and unlock the design of new effective models that benefit from both worlds.
@@ 1 Introduction
Graph Neural Networks (GNNs) sperduti1993encoding ; gori2005new ; scarselli2008graph ; NN4G ; bruna2014spectral ; defferrard2017convolutional have become a widely used architecture for processing information on graph domains. Most GNNs operate via message passing, where information is exchanged between neighboring nodes, giving rise to Message-Passing Neural Networks (MPNNs). Some of the most popular instances of this type of architecture include GCN kipf2017semisupervised , GAT Velickovic2018GraphAN , GIN xu2018powerful , and GraphSAGE hamilton2017inductive .
Despite its widespread use, this paradigm also suffers from some fundamental limitations. Most importantly, we highlight the issue of feature collapse (sometimes termed as over-smoothing) nt2019revisiting ; cai2020note , where feature representations become exponentially similar as the number of layers increases, and over-squashing Alon2021OnTB ; topping2021understanding ; di2023over , which describes the difficulty of propagating information across faraway nodes, as the exponential growth in a node’s receptive field results in many messages being compressed into fixed-size vectors. Although these two issues have been studied extensively, and there exists evidence that they are trade-offs of each other giraldo2023trade , there is no unified theoretical framework that explains why architectures that solve these problems work and whether there exists a common underlying cause that governs these problems.
In this work, we analyze these issues arising from GNN depth from the lens of vanishing gradients. In particular, we ask several questions about the appearance and consequences of this phenomenon in GNNs: (i) How prone are GNNs to gradient vanishing? (ii) What is the effect of gradient vanishing on node feature evolution? (iii) Can preventing vanishing gradients effectively mitigate over-squashing and enable long-range modelling? (iv) Can methods used in the non-linear hochreiter1997long ; pascanu2013difficulty ; beck2024xlstm and more recently, linear gu2023mamba ; orvieto2023resurrecting recurrent neural network (RNN) literature be effective at dealing with feature collapse and long-range modeling? We answer these questions by providing a novel perspective that bridges the gap between recurrent and graph learning.
Figure 1: Latent evolution of 2-dimensional node features when passing through layers of a GNN-SSM with $\mathrm{eig}(\Lambda)\approx 1$. Node states evolve in a norm-preserving manner, without collapsing or contracting. The blue lines indicate how each node feature evolves across layers, i.e., as more layers are added. Each circle corresponds to a node’s 2D feature. Circles connected by a blue line represent the same node across successive layers.The color of each circle encodes the norm of the node feature, and the vector field indicates direction.
@@
  • In Section 3, we explore the connection between GNNs and recurrent models and demonstrate how classical GNNs are susceptible to a phenomenon we term extreme gradient vanishing. We propose GNN-SSM, a GNN model that is written as a state-space model, allowing for better control of the spectrum of the Jacobian.
  • In Section 4, we show how vanishing gradients contribute to node feature evolution, providing a more precise explanation for why GNNs struggle with depth and how node feature collapse emerges via spectral analysis of the layer-wise Jacobians. We show that GNN-SSMs are able to exactly control the rate of collapse of deep representations.
@@ 2.1 Message Passing Neural Networks
for $k=\{1,\ldots,K\}$, where $\psi^{(k)}$ is a permutation-invariant aggregation function and $\phi^{(k)}$ combines the incoming messages from one’s neighbors with the previous embedding of oneself to produce an updated representation. The most commonly used aggregation function takes the form
@@
where $\tilde{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, and $\mathbf{D}\in\mathbb{R}^{n\times n}$ is a diagonal matrix where $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. One can also consider a matrix representation of the features $\mathbf{H}^{(k)}\in\mathbb{R}^{n\times d_{k}}$. Throughout the paper, we will use the terms GNN and MPNN interchangeably, and will generally consider the most widely used instance of GNNs, which are Graph Convolutional Networks (GCNs) (kipf2017semisupervised, ) whose matrix update equation is given by:
@@
where $\hat{\mathbf{A}}=\left(\mathbf{D}+\mathbf{I}\right)^{-1/2}\left(\mathbf{A}+\mathbf{I}\right)\left(\mathbf{D}+\mathbf{I}\right)^{-1/2}$ is the adjacency matrix with added self connections through the identity matrix $\mathbf{I}$, and $\sigma(\cdot)$ is a nonlinearity. Our analysis also applies to Graph Attention Networks (GATs) Velickovic2018GraphAN , where the fixed normalized adjacency is replaced by a learned adjacency matrix which dynamically modulates connectivity while preserving the key spectral properties used in our analysis.
@@ 2.2 Recurrent Neural Networks
This type of approach has deep connections with ideas from dynamical systems strogatz2018nonlinear and chaotic systems engelken2023lyapunov . These ideas have become more relevant in recent work Gu2019LearningMR ; orvieto2023resurrecting , where the nonlinearity in (4) is removed in the interest of parallelization and the ability to directly control the dynamics of the system through the eigenvalues of state transition matrix $\mathbf{W}_{h}$. We note that these types of approaches are also popular in the reservoir computing literature jaeger2001echo , where the state transition matrix is left untrained and more emphasis is placed on the dynamics of the model.
@@ 2.3 The Vanishing and Exploding Gradient Problem
$\frac{\partial\mathcal{L}}{\partial\mathbf{\boldsymbol{\theta}}^{(i)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{(K)}}\Bigg(\prod_{k=i+1}^{K}\frac{\partial\mathbf{H}^{(k)}}{\partial\mathbf{H}^{(k-1)}}\Bigg)\frac{\partial\mathbf{H}^{(i)}}{\partial\mathbf{\boldsymbol{\theta}}^{(i)}},$ (5)
where matrix $\mathbf{H}^{(k)}$ in an RNN will contain a single state vector. As identified by (pascanu2013difficulty, ), a major issue in training this type of models arises from the product Jacobian, given by:
@@
In general, we have that if $\mathbf{J}_{k}||_{2}\approx\lambda$ for all layers then $\mathbf{J}||_{2}\leq\lambda^{K-i}$. This means that we require $\lambda\approx 1$ for gradients to neither explode nor vanish, a condition also known as edge of chaos.
@@ 2.4 Over-smoothing, Over-squashing, and Vanishing Gradients in GNNs
Node Feature Collapse and Over-smoothing.
GNNs are known to not perform well at large depths li2019deepgcns . This issue has been heavily linked with the issue of over-smoothing cai2020note ; oono2020graph , which describes the tendency of GNNs to produce smoother representations as more and more layers are added. In Section 4, we study this issue from the lens of vanishing gradients and show that the performance degradation of GNNs has a much more simple explanation: it occurs due to the norm-contracting nature of GNN updates, which is also intimately related to some notions of smoothing presented in the literature rusch2023survey . [^2] We elaborate on the definition we use in the paper around over-smoothing as it related to the broader literature in Appendix E.
@@
Over-squashing Alon2021OnTB ; topping2021understanding ; di2023over ; barbero2024transformers was originally introduced as a bottleneck resulting from ‘squashing’ into node representations amounts of information that are growing potentially exponentially quickly due to the topology of the graph. It is often characterized by the quantity $\left\lVert\partial\mathbf{h}_{u}^{(K)}/\partial\mathbf{h}_{v}^{(0)}\right\rVert$ being low, implying that the final representation of node $u$ is not very sensitive to the initial representation at some other node $v$. While the relationship between over-squashing and vanishing gradients was hinted at by (di2023over, ), in Section 5 we explore this relationship in detail by showing that techniques aimed to mitigate vanishing gradients in sequence models help to mitigate over-squashing in GNNs.
@@
Vanishing gradients have been extensively studied in RNNs bengio1994learning ; hochreiter1997long ; pascanu2013difficulty , while this problem has been surprisingly mostly overlooked in the GNN community. For a detailed discussion on the relevant literature, we point the reader to the Appendix F. We simply highlight that there are works that have seen success in taking ideas from sequence modelling rusch2022graph ; gravina2022anti ; wang2024mamba ; behrouz2024graphmamba ; kiani2024unitary or signal propagation epping2024graph ; scholkemper2024residual and bridging them to GNNs, but they rarely have a detailed discussion on vanishing gradients. In Section 5, we show that vanishing gradient mitigation techniques from RNNs seem to be very effective towards the mitigation of feature collapse and over-squashing in GNNs and argue that the two communities have very aligned problems and goals.
@@ 3.1 Similarities and differences between learning on sequences and graphs
The GNN architectures that first popularized deep learning on graphs (bruna2014spectral, ; defferrard2017convolutional, ) were initially presented as a generalization of Convolutional Neural Networks (CNNs) to irregular domains. GCNs (kipf2017semisupervised, ) subsequently restricted the architecture in (defferrard2017convolutional, ) to a one-hop neighborhood. While this is still termed “convolutional” (due to weight sharing across nodes), the iterative process of aggregating information from each node’s neighborhood can also be viewed as recurrent-like state updates.
If we consider an RNN unrolled over time, it forms a directed path graph feeding into a state node with a self-connection, making it a special case of a GNN. Conversely, node representations in GNNs can be stacked using matrix vectorization, allowing us to interpret GNN layer operations as iterative state updates. This connection suggests that the main difficulty faced by RNNs, namely the vanishing and exploding gradients problem (pascanu2013difficulty, ), may likewise hinder the learning ability of GNNs. We note, however, that one key difference between RNNs and GNNs is that RNN memory only depends on how much information is dissipated by the model during the hidden state update, whereas GNNs normalize messages by the inverse node degree, which introduces an additional information dissipation step that we will explore in more detail in Section 5.
@@ 3.2 Graph convolutional and attentional models are prone to extreme gradient vanishing
{eqbox}
Lemma 3.1 (Spectrum of layer-wise Jacobian’s singular values).
@@
$\bigl\{\,\lambda_{i}^{2}\,\mu_{j}\;\bigm|\;i=1,\ldots,n,\;\;j=1,\ldots,d_{k}\bigr\}.\\$
{eqbox}
@@
$\vskip-17.07182pt\mathbb{E}\bigl[\gamma_{i,j}\bigr]$ $=\lambda_{i}^{2}\,\sigma^{2},$ (7)
@@
Theorem 3.2 shows that the singular-value spectrum of the Jacobian is modulated by the squared spectrum of the normalized adjacency. Since $|\lambda_{i}|\leq 1$ for all eigenvalues of the normalized adjacency, the ability of GCNs to propagate gradients is in expectation worse than that of RNNs or MLPs. In particular, iterating these operations causes the majority of the spectrum to shrink to zero more quickly than in classical deep linear saxe2013exact or nonlinear pennington2017resurrecting networks. Moreover, using sigmoidal activations and orthogonal weights will not push the singular-value spectrum to the edge of chaos as in (pennington2017resurrecting, ), due to the additional contraction from the adjacency.
Figure 2: Left: Histogram of eigenvalue modulus of the Jacobian for linear, linear convolutional, and nonlinear convolutional layers.  Middle: Vectorized Jacobian for GCN.  Right: Vectorized Jacobian for GCN-SSM with $\mathrm{eig}(\Lambda)\approx 1$, $\mathrm{eig}(B)\approx 0.1$.
The effect of each operation on the layer-wise Jacobian is empirically demonstrated in Figure 2, which also showcases the contraction effect of the normalized adjacency. The figure reveals that even a single layer’s Jacobian exhibits a long tail of squared singular values near zero. This spectral structure leads to ill-conditioned gradient propagation and non-isometric (not norm-preserving) signal dynamics. The same results hold for GATs, as the adjacency still exhibits a contractive spectral structure despite being learned during training.